OpenStudy (anonymous):

Please solve 4x - 6x^3 + 6/12x My mom is making me do summer math questions and she decided to have me do this but I'm stuck...

OpenStudy (loser66):

My mom gives me exactly that problem but there is something after that chain of letter, like = 0 . Since my mom's stuff doesn't look like your mom's one. I don't know how to do with yours.

OpenStudy (anonymous):

I think she wrote it wrong... did you're mom give you something close?

OpenStudy (loser66):

it's hers 4x -6x^3 +6/12x=0 and asks me solve it. it is different from yours, right?

OpenStudy (anonymous):

She said that she "was rushed and chopped it off... sorry." And now im looking at the problem but is it possible to divide 4x by 12x and =0?

OpenStudy (whpalmer4):

Here's how to solve \[4x-6x^3+\frac{6}{12}x=0\] My able assistant will take over at this point :-)

OpenStudy (loser66):

show me more, britanny. my mom says " factor" or something like that

OpenStudy (anonymous):

I got as far as: 4x - 6x^3 + 6 /12 /12 /12 3 - 2^3 + 6

OpenStudy (blurbendy):

is it supposed to be:  or as whpalmer4 showed?